The sum of two numbers is six. The sum of the squares of the two numbers is twenty-six. Find the two numbers.
Answer by stanbon(75887)

The sum of two numbers is six. The sum of the squares of the two numbers is twenty-six. Find the two numbers.
-------------------------
Let one of the numbers be "x".
Then the other numberis "6-x".
----------------------------
EQUATION:
x^2 + (6-x)^2 = 26
x^2 + 36 -12x + x^2 = 26
2x^2 + 10 -12x = 0
x^2 - 6x + 5 = 0
(x-5)(x-1) = 0
x = 5 or x = 1
====================
Cheers,
Stan H.